# An Overview of Systems-Theoretic Guarantees in Data-Driven Model Predictive Control

### v1

@@ 4.1.2 Regularization in data-driven MPC
While robust data-driven MPC (Algorithm 4.1) differs from the nominal approach (Algorithm 3.1) in multiple aspects, the most crucial modification is the regularization of $\mathbf{\alpha}(t)$ in the cost. Indeed, satisfactory practical results can often be obtained even without the slack variable $\mathbf{\sigma}(t)$. On the contrary, the absence of the regularization of $\mathbf{\alpha}(t)$ can pose difficulties already for tiny noise levels or even in the complete absence of noise due to numerical inaccuracies. Intuitively, this regularization can be understood as reducing the complexity of the employed prediction model based on the Fundamental Lemma, analogous to regularization methods in linear regression. In particular, due to the multiplication of $\mathbf{\alpha}(t)$ with the noisy output Hankel matrix, values $\mathbf{\alpha}(t)$ with small norm are desired to reduce the prediction error due to the noise.
Since the introduction of regularization as a heuristic in [12], numerous recent works have illuminated the role of regularization in data-driven MPC. First, we note that regularization plays a crucial role in the closed-loop analysis of data-driven MPC, where it allows to control the size of $\mathbf{\alpha}(t)$, which influences the prediction error [51, 58, 66, 69, 70, 72]. Moreover, a variety of works have focused on the effect of regularization in data-driven finite-horizon optimal control, i.e., applying the full input sequence computed via the optimization problem 8 in open loop. When considering stochastic noise, the regularized optimization problem is equivalent to a min-max optimal control problem subject to distributional uncertainty, i.e., regularization serves as a robustification against noise [73]. The regularization can also be derived through the lens of maximum likelihood estimation [74, 75] or by minimizing the expected cost rather than the nominal cost [76]. Further, [77] has proposed $\gamma$-data-driven predictive control ($\gamma$-DDPC), which decomposes the data-driven optimization problem 8 into two separate steps and avoids the tuning of the regularization parameter $\lambda_{\mathbf{\alpha}}$. The $\gamma$-DDPC scheme also allows to construct regularization terms which systematically reduce the prediction error and can be tuned without running additional experiments [78, 79]. This framework has been generalized in [80] via the Final Control Error, which can be used to construct a data-driven MPC scheme that minimizes the actual prediction error due to the noise and can be applied without tuning regularization parameters. Moreover, additional causality information can be used to improve performance in $\gamma$-DDPC [81]. Notably, equivalence between the classical regularized problem as in Algorithm 4.1 and $\gamma$-DDPC can be established for suitable choices of the regularization parameters [82].
@@ 4.1.3 Closed-loop guarantees for stochastic noise
The stochastic Fundamental Lemma can be used to construct a data-driven MPC scheme for unknown stochastic systems subject to chance constraints. In [91], it is shown that this scheme admits closed-loop guarantees on recursive feasibility and practical stability as well as closed-loop performance bound. Here, closed-loop guarantees are ensured via terminal ingredients, i.e., a terminal cost function and a terminal region constraint, which can be constructed based only on data. Subsequent works have extended the result in [91], which considers state measurements and a binary selection of the initial condition as either the measured state or a backup in case of infeasibility (the prediction from the previous time step). In particular, [92] provides an MPC scheme with analogous guarantees for input-output measurements and [93] optimizes the initial condition, interpolating between the measured state and the backup. Since stochastic data-driven MPC based on PCE admits an increased computational cost, a tailored multiple shooting approach for complexity reduction was derived in [94].
As an alternative to the PCE approach, closed-loop guarantees of data-driven MPC with stochastic noise can be ensured by resorting to techniques from tube-based robust MPC, i.e., computing a constraint tightening offline such that closed-loop guarantees can be established [95]. All previously mentioned approaches for stochastic data-driven MPC require access to past disturbance realizations and, therefore, they derive data-driven estimation procedures. As an alternative, the recent work [96] constructs an explicit data-driven parametrization of consistent disturbance realizations from which new samples can be drawn, allowing to set up a sampling-based stochastic data-driven MPC scheme which guarantees closed-loop properties via a robust first-step constraint.
@@ 4.2 Data-driven MPC for more advanced control objectives
Further data-driven MPC formulations in the literature include explicit MPC [100], which relies on an explicit solution of the underlying quadratic program via the Karush-Kuhn-Tucker conditions, and economic MPC [101], which addresses stage cost functions that are not positive (semi-)definite and can also handle unknown linear cost functions. In the context of networked and cyber-physical control systems, data-driven MPC schemes have been proposed that address resilience against denial-of-service attacks [102] as well as self-triggered [103] and event-triggered [104] MPC formulations. Further, various contributions have been made on safe data-driven control via the Fundamental Lemma. This includes the design of a data-driven safety filter which equips any controller with safety guarantees by ensuring that unsafe regions are avoided [105, 106], the combination of a data-driven MPC scheme with a funnel controller [107], and enhancing data-driven MPC with control barrier certificates [108]. The Fundamental Lemma as well as data-driven state-feedback design approaches [109, 55] have also inspired alternative data-driven MPC schemes relying on linear matrix inequalities [110, 111]. These approaches consider an infinite-horizon cost and a state-feedback parametrization of the input, and they may be less conservative than data-driven MPC when guaranteeing constraint satisfaction in the presence of noise, see [111] for a numerical comparison. Further formulations include data-driven MPC for linear descriptor systems [112] as well as data-driven control for iterative tasks [113]. Finally, while all the above approaches use the Fundamental Lemma for data-driven control, the dual problem of observer design has also been studied: In [114], a data-driven moving horizon estimation scheme is developed which proves practical exponential stability of the estimator in case of noisy data.
In conclusion, the literature contains a variety of data-driven MPC schemes for different problem setups, control objectives, and system classes. While many of these approaches are inspired by analoguous developments in model-based MPC, the Fundamental Lemma allows to set up MPC schemes for controlling unknown systems based only on measured data, which brings unique challenges for the theoretical analysis, in particular in the presence of uncertainty (e.g., noisy data or disturbances).
@@ 5 DATA-DRIVEN MPC FOR NONLINEAR SYSTEMS
In the literature, two complementary approaches for nonlinear data-driven MPC based on the Fundamental Lemma have been developed, which we discuss in Sections 5.1 and 5.2, respectively. The first class of methods rely on nonlinear versions of the Fundamental Lemma, which can be developed in a tailored fashion for specific system classes by assuming some structural knowledge, e.g., basis functions, and exploiting global linearity in higher-dimensional coordinates. The second class follows an adaptive approach and updates the data used in the Fundamental Lemma online, exploiting that nonlinear systems are locally well-approximated by linear dynamics.
@@ 5.1 Exploiting global linearity via a nonlinear Fundamental Lemma
Each of these nonlinear versions of the Fundamental Lemma allows to design a data-driven MPC scheme for systems in the associated system class. In the following, we show how this can be done for Hammerstein systems via the corresponding Fundamental Lemma from [115]. Analogous MPC schemes can be designed for other nonlinear system classes when using the corresponding version of the Fundamental Lemma, cf. above. A Hammerstein system is a series interconnection of a static nonlinearity and an LTI system, i.e.,
@@
compare [115, Proposition 5]. This result exploits the linearity of the system which maps the auxiliary input $\mathbf{v}$, whose components involve the actual input $\mathbf{u}$ and the known basis functions $\psi_{i}$, to the output $\mathbf{y}$. Since Equation 11 allows to accurately predict input-output trajectories of the given Hammerstein system, it can be used to set up a data-driven MPC scheme precisely as in the linear noise-free case considered in Section 3.1. Again, closed-loop guarantees can be established via appropriate modifications of the basic optimization problem, e.g., terminal equality constraints, terminal cost and terminal region constraints, or a sufficiently long prediction horizon. The main limitation of this approach, beyond its restriction to Hammerstein systems and knowledge of the $\psi_{i}$’s, is that the resulting optimization problem to be solved at every time step is typically non-convex due to the nonlinear coupling of the decision variables $\mathbf{\bar{u}}$ and $\mathbf{\bar{v}}$ in Equation 12.
@@ 5.2 Exploiting local linearity via online data updates
As an alternative to nonlinear versions of the Fundamental Lemma, one can resort to a linear Fundamental Lemma and update the data used for prediction in the Hankel matrices online. Since (smooth) nonlinear systems are locally well-approximated by linear dynamics, the resulting predictions are guaranteed to be accurate under suitable conditions. In the following, we discuss this idea in more detail and introduce the data-driven MPC approach from [70], which provides closed-loop guarantees when controlling unknown nonlinear systems based only on input-output data.
@@
with Jacobians and remainder terms
$\mathbf{A}_{\mathbf{x}_{t}}=\frac{\partial\mathbf{f}}{\partial\mathbf{x}}\Big{\lvert}_{\mathbf{x}_{t}},\>\>\mathbf{C}_{\mathbf{x}_{t}}=\frac{\partial\mathbf{h}}{\partial\mathbf{x}}\Big{\lvert}_{\mathbf{x}_{t}},\>\>\mathbf{e}_{\mathbf{x}_{t}}=\mathbf{f}(\mathbf{x}_{t})-\mathbf{A}_{\mathbf{x}_{t}}\mathbf{x}_{t},\>\>\mathbf{r}_{\mathbf{x}_{t}}=\mathbf{h}(\mathbf{x}_{t})-\mathbf{C}_{\mathbf{x}_{t}}\mathbf{x}_{t}.$ (15)
Before introducing the data-driven MPC scheme, we take a slight detour to model-based MPC. The work [131] proposes an MPC scheme to control nonlinear systems based on the linearized dynamics in Equation 14, i.e., an MPC scheme based on precise model knowledge. The approach includes a terminal equality constraint w.r.t. an artificial equilibrium $\mathbf{x}^{\mathrm{s}}(t)$ for the linearized dynamics, i.e., a vector $\mathbf{x}^{\mathrm{s}}(t)$ satisfying
@@
for some equilibrium input $\mathbf{u}^{\mathrm{s}}(t)$ and output $\mathbf{y}^{\mathrm{s}}(t)$, compare [61]. The distance of the artificial equilibrium $\mathbf{x}^{\mathrm{s}}(t)$ to the actual setpoint is penalized in the cost. If the corresponding cost weight is sufficiently small, then $\mathbf{x}^{\mathrm{s}}(t)$ is encouraged to remain close to the initial state $\mathbf{x}_{t}$. If $\mathbf{x}_{t}$ is, in turn, close to the steady-state manifold of the nonlinear system, then the full predicted trajectory $\mathbf{\bar{x}}(t)$ remains in a small region around $\mathbf{x}_{t}$, i.e., in a region where the linearized dynamics in Equation 14 provide an accurate prediction. Based on this idea, it can be shown that the MPC scheme based on the linearized dynamics indeed exponentially stabilizes the setpoint in closed loop. Notably, this guarantee does not only hold locally around the setpoint, but the region of attraction is a neighborhood of the entire steady-state manifold, which can be significantly larger. Roughly speaking, e.g., in an autonomous driving application, successful tracking of a position can be guaranteed when starting from any initial position, under the restriction that the car does not drive too fast, i.e., it remains close to the steady-state manifold.
@@
By using an MPC scheme with artificial equilibrium as in [131], it can be enforced that the system does not change too rapidly in closed loop and, therefore, the trajectory $\{\mathbf{u}_{k},\mathbf{y}_{k}\}_{k=t-N}^{t-1}$ of the nonlinear system is close to a trajectory from the dynamics linearized at $\mathbf{x}_{t}$. More precisely, if a certain cost matrix in the MPC optimization problem is chosen small enough, then the closed-loop trajectory moves slowly such that Equation 17 provides locally accurate predictions for the nonlinear system. Combining these ideas in a rigorous fashion, it can be shown that the setpoint is practically exponentially stable for the closed-loop system, i.e., the trajectory converges to a neighborhood around the setpoint [70, Theorem 2]. The size of this neighborhood depends on the distance between data samples generated initially at time $t=0,\dots,N-1$, e.g., through random exploration. As in the model-based MPC scheme from [131], the guaranteed region of attraction of the data-driven MPC scheme is a neighborhood of the steady-state manifold of the underlying system.
There are several notable differences of this approach to those based on nonlinear versions of the Fundamental Lemma explained in Section 5.1. On the theoretical side, data-driven MPC with online data updates provides guarantees for system classes which are complementary to those considered in Section 5.1, relying on local arguments and therefore requiring entirely different and partially less restrictive assumptions: not restricted to systems which can be represented linearly in suitable known coordinates; input-affine dynamics (can be enforced via an input transformation); smooth dynamics; properties that are also needed in model-based MPC, e.g., controllability, observability. On the practical side, data-driven MPC with online data updates only requires input-output measurements for its implementation but no choice of basis functions for nonlinear components as in most approaches from Section 5.1. Further, since the approach relies on the affine Fundamental Lemma, the resulting optimization problems are strictly convex quadratic programs, in contrast to data-driven MPC based on nonlinear basis functions. Further, data-driven MPC with online data updates has been successfully applied in real-world experiments, including a four-tank system [14] and a complex soft robot [17]. A related data-driven MPC scheme which also uses online data updates but does not admit closed-loop guarantees was applied to a building control experiment in [26].

### v2

@@ 4.1.2 Regularization in data-driven MPC
While robust data-driven MPC (Algorithm 4.1) differs from the nominal approach (Algorithm 3.1) in multiple aspects, the most crucial modification is the regularization of $\mathbf{\alpha}(t)$ in the cost. Indeed, satisfactory practical results can often be obtained even without the slack variable $\mathbf{\sigma}(t)$. On the contrary, the absence of the regularization of $\mathbf{\alpha}(t)$ can pose difficulties already for low noise levels or even in the complete absence of noise due to numerical inaccuracies. Intuitively, this regularization can be understood as reducing the complexity of the employed prediction model based on the Fundamental Lemma, analogous to regularization methods in linear regression. In particular, due to the multiplication of $\mathbf{\alpha}(t)$ with the noisy output Hankel matrix, values $\mathbf{\alpha}(t)$ with small norm are desired to reduce the prediction error due to the noise.
Since the introduction of regularization as a heuristic in [12], numerous recent works have illuminated the role of regularization in data-driven MPC. First, we note that regularization plays a crucial role in the closed-loop analysis of data-driven MPC, where it allows one to control the size of $\mathbf{\alpha}(t)$, which influences the prediction error [51, 58, 68, 71, 72, 67]. Moreover, a variety of works have focused on the effect of regularization in data-driven finite-horizon optimal control, i.e., applying the full input sequence computed via the optimization problem 8 in open loop. When considering stochastic noise, the regularized optimization problem is equivalent to a min-max optimal control problem subject to distributional uncertainty, i.e., regularization serves as a robustification against noise [74]. The regularization can also be derived through the lens of maximum likelihood estimation [75, 76] or by minimizing the expected cost rather than the nominal cost [77]. Further, [78] has proposed $\gamma$-data-driven predictive control ($\gamma$-DDPC), which decomposes the data-driven optimization problem 8 into two separate steps and avoids the tuning of the regularization parameter $\lambda_{\mathbf{\alpha}}$. The $\gamma$-DDPC scheme also allows one to construct regularization terms which systematically reduce the prediction error and can be tuned without running additional experiments [79, 80]. This framework has been generalized in [81] via the Final Control Error, which can be used to construct a data-driven MPC scheme that minimizes the actual prediction error due to the noise and can be applied without tuning regularization parameters. Moreover, additional causality information can be used to improve performance in $\gamma$-DDPC [82]. Notably, equivalence between the classical regularized problem as in Algorithm 4.1 and $\gamma$-DDPC can be established for suitable choices of the regularization parameters [83].
@@ 4.1.3 Closed-loop guarantees for stochastic noise
The stochastic Fundamental Lemma can be used to construct a data-driven MPC scheme for unknown stochastic systems subject to chance constraints on the input and state. In [92], it is shown that this scheme admits closed-loop guarantees on recursive feasibility and practical stability as well as closed-loop performance bound. Here, closed-loop guarantees are ensured via terminal ingredients, i.e., a terminal cost function and a terminal region constraint, which can be constructed based only on data. Subsequent works have extended the result in [92], which considers state measurements and a binary selection of the initial condition as either the measured state or a backup in case of infeasibility (the prediction from the previous time step). In particular, [93] provides an MPC scheme with analogous guarantees for input-output measurements and [94] optimizes the initial condition, interpolating between the measured state and the backup. Since stochastic data-driven MPC based on PCE admits an increased computational cost, a tailored multiple shooting approach for complexity reduction was derived in [95].
As an alternative to the PCE approach, closed-loop guarantees of data-driven MPC with bounded stochastic noise can be ensured by resorting to techniques from tube-based robust MPC, i.e., computing a constraint tightening offline such that closed-loop guarantees can be established [96]. All previously mentioned approaches for stochastic data-driven MPC require access to past disturbance realizations and, therefore, they derive data-driven estimation procedures. As an alternative, the recent work [97] constructs an explicit data-driven parametrization of consistent disturbance realizations from which new samples can be drawn, leading to a sampling-based stochastic data-driven MPC scheme which guarantees closed-loop properties via a robust first-step constraint.
@@ 4.2 Data-driven MPC for more advanced control objectives
Further data-driven MPC formulations in the literature include explicit MPC [101], which relies on an explicit solution of the underlying quadratic program via the Karush-Kuhn-Tucker conditions, and economic MPC [102], which addresses stage cost functions that are not positive (semi-)definite and can also handle unknown linear cost functions. In the context of networked and cyber-physical control systems, approaches have been proposed that address resilience against denial-of-service attacks [103] as well as self-triggered [104] and event-triggered [105] MPC formulations. Further, various contributions have been made on safe data-driven control via the Fundamental Lemma. This includes the design of a data-driven safety filter which equips any controller with safety guarantees [106, 107], the combination of a data-driven MPC scheme with a funnel controller [108], and enhancing data-driven MPC with control barrier certificates [109]. The Fundamental Lemma as well as data-driven state-feedback design approaches [55, 110] have also inspired alternative data-driven MPC schemes relying on linear matrix inequalities [111, 112]. These approaches consider an infinite-horizon cost and a state-feedback parametrization of the input, and they may be less conservative than data-driven MPC when guaranteeing constraint satisfaction in the presence of noise, see [112] for a numerical comparison. Further formulations include an extended Kalman Filter for reducing computational complexity [113], data-driven MPC for linear descriptor systems [114], and data-driven control for iterative tasks [115]. Finally, while all the above approaches use the Fundamental Lemma for data-driven control, the dual problem of observer design has also been studied: In [116], a data-driven moving horizon estimation scheme is developed which guarantees practical exponential stability of the estimator in case of noisy data.
In conclusion, the literature contains a variety of data-driven MPC schemes for different problem setups, control objectives, and system classes. While many of these approaches are inspired by analoguous developments in model-based MPC, the Fundamental Lemma can be used to set up MPC schemes for controlling unknown systems based only on measured data, which brings unique challenges for the theoretical analysis, in particular in the presence of uncertainty (e.g., noisy data or disturbances).
@@ 5 DATA-DRIVEN MPC FOR NONLINEAR SYSTEMS
In the literature, two complementary approaches for nonlinear data-driven MPC based on the Fundamental Lemma have been developed, which we discuss in Sections 5.1 and 5.2, respectively. The first class of methods rely on nonlinear versions of the Fundamental Lemma, which can be developed in a tailored fashion for specific system classes by assuming some structural knowledge, e.g., basis functions, and exploiting global linearity in higher-dimensional coordinates. The second class follows an adaptive approach and updates the data used in the Fundamental Lemma online, exploiting that nonlinear systems are locally well-approximated by affine dynamics. Throughout this section, we consider noise-free data but we note that the majority of the approaches can handle noisy data with modifications analogous to those discussed in Section 4.1.
@@ 5.1 Exploiting global linearity via a nonlinear Fundamental Lemma
Each of these nonlinear versions of the Fundamental Lemma allows one to design a data-driven MPC scheme for systems in the associated system class. In the following, we show how this can be done for Hammerstein systems via the nonlinear Fundamental Lemma from [117]. Analogous MPC schemes can be designed for other nonlinear system classes when using the corresponding version of the Fundamental Lemma. A Hammerstein system is a series interconnection of a static nonlinearity and an LTI system, i.e.,
@@
compare [117, Proposition 5]. This result exploits the linearity of the system which maps the auxiliary input $\mathbf{v}$, whose components involve the actual input $\mathbf{u}$ and the known basis functions $\psi_{i}$, to the output $\mathbf{y}$. Since Equation 11 accurately predicts input-output trajectories of the given Hammerstein system, it can be used to set up a data-driven MPC scheme precisely as in the linear noise-free case considered in Section 3.1. Again, closed-loop guarantees can be established via appropriate modifications of the basic optimization problem, e.g., terminal equality constraints, terminal cost and terminal region constraints, or a sufficiently long prediction horizon. The main limitation of this approach, beyond its restriction to Hammerstein systems and knowledge of the $\psi_{i}$’s, is that the resulting optimization problem to be solved at every time step is typically non-convex due to the nonlinear coupling of the decision variables $\mathbf{\bar{u}}$ and $\mathbf{\bar{v}}$ in Equation 12.
@@ 5.2 Exploiting local linearity via online data updates
Alternatively, one can resort to an affine Fundamental Lemma and update the data used for prediction in the Hankel matrices online. Since (smooth) nonlinear systems are locally well-approximated by affine dynamics, the resulting predictions are guaranteed to be accurate under suitable conditions. In the following, we discuss this idea in more detail and introduce the data-driven MPC approach from [72], which provides closed-loop guarantees when controlling unknown nonlinear systems based only on input-output data.
@@
with Jacobians $\mathbf{A}_{\mathbf{x}_{t}}=\frac{\partial\mathbf{f}}{\partial\mathbf{x}}\Big\lvert_{\mathbf{x}_{t}}$, $\mathbf{C}_{\mathbf{x}_{t}}=\frac{\partial\mathbf{h}}{\partial\mathbf{x}}\Big\lvert_{\mathbf{x}_{t}}$ and remainder terms $\mathbf{e}_{\mathbf{x}_{t}}=\mathbf{f}(\mathbf{x}_{t})-\mathbf{A}_{\mathbf{x}_{t}}\mathbf{x}_{t}$, $\mathbf{r}_{\mathbf{x}_{t}}=\mathbf{h}(\mathbf{x}_{t})-\mathbf{C}_{\mathbf{x}_{t}}\mathbf{x}_{t}$. Before introducing the data-driven MPC scheme, we take a slight detour to model-based MPC. The work [133] proposes an MPC scheme to control nonlinear systems based on the linearized dynamics in Equation 14, i.e., an MPC scheme based on precise model knowledge. The approach includes a terminal equality constraint with respect to an artificial equilibrium $\mathbf{x}^{\mathrm{s}}(t)$ for the linearized dynamics, i.e., a vector $\mathbf{x}^{\mathrm{s}}(t)$ satisfying
@@
for some equilibrium input $\mathbf{u}^{\mathrm{s}}(t)$ and output $\mathbf{y}^{\mathrm{s}}(t)$, compare [61]. The distance of the artificial equilibrium $\mathbf{x}^{\mathrm{s}}(t)$ to the actual setpoint is penalized in the cost. If the corresponding cost weight is sufficiently small, then $\mathbf{x}^{\mathrm{s}}(t)$ is encouraged to remain close to the initial state $\mathbf{x}_{t}$. If $\mathbf{x}_{t}$ is, in turn, close to the steady-state manifold of the nonlinear system, then the full predicted trajectory $\mathbf{\bar{x}}(t)$ remains in a small region around $\mathbf{x}_{t}$, where the linearized dynamics in Equation 14 provide an accurate prediction. Based on this idea, it can be shown that the MPC scheme using the linearized dynamics exponentially stabilizes the setpoint. Notably, this guarantee does not only hold locally around the setpoint, but the region of attraction is a neighborhood of the entire steady-state manifold, which can be significantly larger. Roughly speaking, e.g., in an autonomous driving application, successful tracking of a position can be guaranteed when starting from any initial position, under the restriction that the car does not drive too fast, i.e., it remains close to the steady-state manifold.
@@
By using an MPC scheme with artificial equilibrium as in [133], it can be enforced that the system does not change too rapidly in closed loop and, therefore, the trajectory $\{\mathbf{u}_{k},\mathbf{y}_{k}\}_{k=t-N}^{t-1}$ of the nonlinear system is close to a trajectory from the dynamics linearized at $\mathbf{x}_{t}$. More precisely, if a certain cost matrix in the MPC optimization problem is chosen small enough, then the closed-loop trajectory moves slowly such that Equation 16 provides locally accurate predictions for the nonlinear system. Combining these ideas in a rigorous fashion, it can be shown that the setpoint is practically exponentially stable for the closed-loop system, i.e., the trajectory converges to a neighborhood around the setpoint [72, Theorem 2]. The size of this neighborhood depends on the distance between data samples generated initially at time $t=0,\dots,N-1$, e.g., through random exploration. As in the model-based MPC scheme from [133], the guaranteed region of attraction of the data-driven MPC scheme is a neighborhood of the steady-state manifold of the underlying system.
There are several notable differences of this approach to those based on nonlinear versions of the Fundamental Lemma explained in Section 5.1. On the theoretical side, data-driven MPC with online data updates relies on local arguments and therefore requires different and partially less restrictive assumptions: not restricted to systems which can be represented linearly in suitable known coordinates; input-affine dynamics (can be enforced via an input transformation); smooth dynamics; properties that are also needed in model-based MPC, e.g., controllability, observability. On the practical side, data-driven MPC with online data updates only requires input-output measurements for its implementation but no choice of basis functions for nonlinear components as in Section 5.1. Further, since the approach relies on the affine Fundamental Lemma, the resulting optimization problems are strictly convex quadratic programs, in contrast to data-driven MPC based on nonlinear basis functions. Moreover, data-driven MPC with online data updates has been successfully applied in real-world experiments, including a four-tank system [14] and a complex soft robot [17]. A related data-driven MPC scheme which also uses online data updates but does not admit closed-loop guarantees was applied to a building control experiment in [26].